OpenStudy (anonymous):

I have 3 questions and I have a pretty good idea of an answer for some of them but im not 100% sure if they are right so ill see if I get the same answer: 1. If bh=42 what is de 2.If ab=36 what is jf 3.If ab=36 what is ke

OpenStudy (anonymous):

Do you have a diagram to go with this question?

OpenStudy (anonymous):

Yea I posted an attachment ill repost it

OpenStudy (anonymous):

thanks - didn't see it before

OpenStudy (anonymous):

I think it was my fault so dont worry hahah

OpenStudy (anonymous):

looks like you've got 6 congruent segments across the bottom with a total length of 42, so what do you think DE equals?

OpenStudy (anonymous):

I believe de = 7 but Its important that I get it right so thats why I posted this

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Now for the others. You have a triangle with one side of 36 (that's AB) and another side of 42 (BH). All the other triangles as you move across from left to right are similar to the biggest triangle (ABH.) Are you with me?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Since all the triangles are similar, the proportions of their sides are the same. So 36/42, which reduces to 6/7, is a common proportion in all of the triangles. For example, triangle JFH

OpenStudy (anonymous):

has a base of 14 (7*2) so to find jf, set up a proportion of 6/7=x/14 to get JF.

OpenStudy (anonymous):

so JF would be?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

good! Now do the same thing for KE. Figure the length of the base of triangle KEH and solve the proportion

OpenStudy (anonymous):

you doing OK?

OpenStudy (anonymous):

Yea I got 18 but im not really sure if its right

OpenStudy (anonymous):

that's what I got as well: x/21=6/7. x=18 I think you've got it!

OpenStudy (anonymous):

Alright well thank you so much!

OpenStudy (anonymous):

my pleasure!